OpenStudy (anonymous):

What is the slope of the line passing through points A and B as shown on the graph below. GRAPH PLOTS; A(−3,8) and B(3,6). Don´t just tell me the answer, please explain thoroughly.

OpenStudy (anonymous):

(y2-y1)/(x2-x1) = (6-8)/(3- -3) = -2/6=-1/3

OpenStudy (anonymous):

A(x1,y1), B(x2,y2)

OpenStudy (anonymous):

get the gradient of the slope as mrj has done 8-6/-3-3 2/-6 -1/3

OpenStudy (anonymous):

thanks